# A Fast, Analytic Empirical Model of the Gaia Data Release 3 Astrometric Orbit Catalog Selection Function

### v1

@@ Abstract
In June 2022, the Gaia mission released a catalog of astrometric orbital solutions for 168,065 binary systems, by far the largest such catalog to date. Unlike previous binary stars catalogs, which were heterogeneous collections of orbits from different surveys and instruments, these orbits were derived using Gaia data alone. Despite this homogeneity, the selection function is difficult to characterize because of choices made in the construction of the catalog. Understanding the catalog’s selection function is required to model and interpret its contents. We use a combination of analytic and empirical prescriptions to construct a function that computes the probability that a binary with a given set of properties would have been published in the Gaia Data Release 3 astrometric orbit catalog. We also construct a binary population synthesis model based on Moe & Di Stefano (2017) to validate our characterization of the selection function, finding good agreement with the actual Gaia NSS catalog, with the exception of the orbital eccentricity distribution. The NSS catalog suggests high-eccentricity orbits are relatively uncommon at intermediate periods $100\lesssim P_{orb}\lesssim 1000$ days. As an example application of the selection function, we estimate the Gaia DR3 detection probabilities of the star + BH binaries Gaia BH1, BH2, and BH3. We also estimate the population of Sun-like star + BH binaries in the Galaxy to be $\sim 5000$ for $400100<P_{orb}<$ 400 day, $\lesssim 2,000$ for $1000400<P_{orb}<$ 1000 day, and $\lesssim 20,000$ for $20001000<P_{orb}<$ 2000 days.
facilities: Gaiasoftware: isochrones (Morton, 2015), galaxia (Sharma et al., 2011), astropy (Astropy Collaboration et al., 2013, 2018), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2019), healpy (Zonca et al., 2019), mwdust (Bovy et al., 2019), COSMIC (Breivik et al., 2020)
@@ 1 Introduction
Gaia is an astrometric space mission surveying over 2 billion stars in the Milky Way (Gaia Collaboration et al., 2016a). Its main mission operated from July 2014 through December 2019, and its extended mission is planned to operate through January 2025. The data products and catalogs generated from the photometry, astrometry, and spectroscopy from Gaia are shared publicly in data releases that occur every several years.
@@
In this paper, we perform an empirical forward model of the Gaia NSS astrometric orbits catalog. We focus on the solutions that have an astrometric orbit (as compared to those that only have a spectroscopic orbit) as our primary goal is to understand the selection function of stellar-mass black hole + star binaries, which have thus far only been found in the astrometric orbits catalog. Our forward modeling approach uses analytic and simple prescriptions to approximate the steps in the construction of the catalog. This is a complementary approach to that of El-Badry et al. (2024), who forward model individual observations at the expense of computational cost; more discussion on the tradeoffs between the two approaches is presented in §6.5.
The rest of this paper is organized as follows. In §2 we review the processing pipeline of the Gaia NSS astrometric orbit catalog and identify the most important steps to model. In §3 we describe the binary population synthesis model. In §4 we forward model the NSS astrometric processing cascade, and the results are presented in §5. Applications, shortcomings, possible extensions of the forward model, and comparison to other work are discussed in §6. We provide a final summary and conclusions in §7.
@@ 2 Gaia astrometric non-single star processing and catalog
The astrometric NSS processing and resultant catalog were designed with the intention of prioritizing accurate binary parameters over completeness. This resulted in a series of many filtering steps to remove potentially spurious solutions, at the expense of many accurate solutions. Our goal is to understand the selection effects of this filtering process. The astrometric NSS processing described in Halbwachs et al. (2023) is summarized below, and the portions most relevant to the forward modeling are highlighted; see also Figure 1 of Halbwachs et al. (2023) for a summary.
@@ 2.1 Astrometric orbit solutions
The Gaia DR3 NSS catalog is divided into four tables depending on the type of NSS model used; each table is further subdivided by the particular sub-model used (see §2.1. “Table contents” of Gaia Collaboration et al. (2023b)). This paper is only concerned with modeling the subset of solutions that have astrometric orbits, i.e. Orbital and AstroSpectroSB1 solutions from the the nss_two_body_orbit table. We collectively refer to these as the NSS astrometric orbit solutions.
@@ 2.2 Selection of stars to be processed
An initial filtering was performed on the entire Gaia DR3 gaia_source  catalog to define the input catalog, a subset of sources that were good NSS candidates. Starting from $1.5\times 10^{9}$ sources with single-star astrometric solutions, several filtering steps were implemented to produce the input catalog of $4.1\times 10^{6}$ stars. Only stars in this input catalog were eligible to be fit with an astrometric NSS model. As described in Halbwachs et al. (2023) §2.1, the filters are as follows:
[…]
  • $\texttt{ruwe}>1.4$ (square root of the reduced $\chi^{2}$, with an empirical renormalization based on magnitude and color of source),
[…]
  • $C^{*}<1.645\sigma_{C^{*}}$ (color excess (ratio between the total flux in BP and RP, and the G-band flux (BP + RP)/G) after correcting for passbands, possibly from a contaminating neighbor).
We only model the ruwe  filter. The ipd_frac_multi_peak, ipd_gof_harmonic_amplitude, and $C^{*}$ filters are intended to remove marginally resolved wide binaries. Such binaries tend to have orbital periods $P_{orb}\gtrsim$ decades, which are much too long to receive an astrometric orbital solution; they have $\texttt{ruwe}>1.4$ because their marginally resolved nature leads to spurious epoch astrometry. The filters on apparent magnitude and visibility periods are accounted for in the $\texttt{ruwe}>1.4$ filter. As ruwe  is a measurement of the significance of deviations from a single-star model, sources that are faint or have too few visibility periods will have astrometric uncertainties too large to have a significant deviation.
2.3 Model cascade
Subsequently, the $4.1\times 10^{6}$ stars from the input catalog were fit with a variety of astrometric NSS models. The models were tried in the following order:
@@ 2.3 Model cascade
  • orbital (Orbital solution, nss_two_body_orbit table; 12 parameters11110 parameters were used for a circular or psuedo-circular orbit (see §5.2 “Pseudo-circular and circular orbital solutions” of Halbwachs et al. (2023)). 15 parameters were used if the orbital solution was combined with a spectroscopic solution (AstroSpectroSB1 solution). The astrometry-only and spectroscopic-only inferred parameters were only combined if they matched (see §5.3. “Selection of the orbital solutions” of Halbwachs et al. (2023)).).
If an acceptable solution was found, subsequent models were not tried. Note that the acceleration models precede the orbital model. Thus, if an acceptable acceleration model was found, an orbital model was not tried, even if an orbital solution would have led to a better fit. We discuss the implications of this choice in §4.3.
@@ 2.4 Orbital solution model acceptance
For sources where an orbital model fit was attempted, there were several conditions determining whether the solution was accepted and included in the final orbital solution catalog. A stringent filter was applied on the significance of the parallax solution ($\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$) where $\varpi$ is the parallax, $\sigma_{\varpi}$ is the parallax error, and $P_{orb}$ is the orbital period. This filter was designed to remove false positives with spurious astrometric orbits and orbital periods related to the Gaia scanning law (Holl et al., 2023), but came at the expense of many valid solutions.
There were also several other less stringent filters. A filter was applied on the significance of the astrometric orbit $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{day}})$ where $a_{0}$ and $\sigma_{a_{0}}$ are the photocenter orbit semimajor axis and its uncertainty. Another filter was applied on the error of the orbital eccentricity $\sigma_{e}<0.079\ln(P_{orb}/\textrm{day})-0.244$. Finally, a filter was also applied on the goodness of fit of the binary solution $F_{2}=\sqrt{9\nu/2}[(\chi^{2}/\nu)^{1/3}+(2/9\nu)-1]<25$, where $\chi^{2}$ is the sum of the squared normalized residuals and $\nu$ is the number of degrees of freedom.
We only model the filters on $\varpi/\sigma_{\varpi}$ and $a_{0}/\sigma_{a_{0}}$. The filter on $\sigma_{e}$ only removes a small percentage of solutions. As described in §5.3 “Selection of the orbital solutions” of Halbwachs et al. (2023), approximately 180,000 orbital solutions remained after applying the filters $F_{2}<25$, $a_{0}/\sigma_{a_{0}}>5$ and $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$; subsequently applying the filters on $\sigma_{e}<0.079\ln(P_{orb}/\textrm{day})-0.244$ and $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{day}}$ only removed an additional 7% of solutions. Thus, the contribution of the $\sigma_{e}$ filter cannot be more than 10% of the final NSS astrometric orbit catalog [^2] Using the modeling of El-Badry et al. (2024), we find that the $\sigma_{e}$ filter only removes 2% of sources, which is 1/3 of the 7% removed by the $a_{0}/\sigma_{a_{0}}$ filter.; this is negligible given the other uncertainties in the forward model. Visual inspection of Figure 3c of Halbwachs et al. (2023) also suggests the $\sigma_{e}$ filter has minimal impact on the final constructed catalog.
We also do not model the filtering on $F_{2}$ because we assume that good orbital solutions will have good fits. $F_{2}$ is expected to follow a normal distribution with mean 0 and standard deviation 1; even if the uncertainties are poorly characterized, $F_{2}<25$ is still a 25-$\sigma$ threshold. It may also allow false positives to pass through, but is unlikely to remove real solutions that have passed the other filtering steps (with the exception of binaries in hierarchical triple systems).
@@ 2.5 Final NSS astrometric orbit catalog
The black outline in Figure 7 shows the distribution of orbital period, parallax, eccentricity, photocenter orbit semimajor axis, renormalized unit weight error, apparent magnitude, and parallax over parallax error for the final set of 168,065 NSS astrometric orbit solutions. Some of the correlations are clearly due to the filtering procedure (e.g. orbital period vs. parallax over parallax error.)
@@ 3.2 Stellar population model (binary properties and evolution, IMF, synthetic photometry)
To forward model the NSS catalog, we cannot directly use the stellar population generated by galaxia since it consists of only single stars. Hence, we combine galaxia positions, metallicities, and ages with a binary star population based on Moe & Di Stefano (2017), which is a joint probability distribution in primary mass $M_{1}$, mass ratio $q=M_{2}/M_{1}$, orbital period, and eccentricity. Moe & Di Stefano (2017) infer the distribution bover $0.8<M_{1}<40M_{\odot}$, $10.1<q<$ 1, $80.15<\log P_{orb}/\textrm{day}<$ 8, and $10<e<$ 1. For reference, about 60% of Sun-like stars are single, while this decreases to nearly 0% for OB-type stars.
@@
We assume that the mass distribution of primaries and singles are drawn from a Kroupa initial mass function (Kroupa, 2001). This results in a multiplicity frequency [^4] Following the definition of Reipurth & Zinnecker (1993), the number of binary systems divided by the number of binary systems and single stars. of 0.21 and a companion probability [^5] Again following Reipurth & Zinnecker (1993), the number of stars in binary systems divided by the total number of stars. of 0.35. This multiplicity frequency might seem significantly lower than binary frequency of $30-40\%$ typically quoted for solar-type stars (e.g., Raghavan et al., 2010; Moe & Di Stefano, 2017). This is because the binary statistics in Moe & Di Stefano (2017) only extend down to $0.8M_{\odot}$; to extrapolate the multiplicity statistics down to $0.08M_{\odot}$, COSMIC assumes the binary fraction decreases linearly from $\approx 40$% at $0.8M_{\odot}$ down to 0 at $0.08M_{\odot}$. This choice of implementation is somewhat arbitrary (given that M-dwarfs appear to have a binary frequency $\sim 20\%$; e.g., Laithwaite & Warren 2020), but ultimately these sources are too faint to be included in the astrometric orbits catalog, so their assumed binary fraction does not matter.
Assuming that the majority of binary systems are unresolved, one galaxia star corresponds to one COSMIC stellar system. This corresponds to 150 million binary systems within 2 kpc at the present day. In order to properly generate the birth population, the stellar progenitors of the compact objects must be accounted for [^6] Note that compact objects are not included in galaxia.. We use the MIST stellar evolution models (Dotter, 2016; Choi et al., 2016) to evolve a population of stars to the same age distribution of the thin disk, thick disk, stellar halo, and bulge. We then determine how many of the primary stars have or will imminently become compact objects (defined as carbon burning for massive stars, and the end of the post-AGB phase for low/intermediate mass stars), and apply a correction factor to the number of stars that need to be generated. This correction is 7.5% for the 5–7 Gyr old population in the thin disk, 11.5% for the 11 Gyr old population in the thick disk, and 12.5% for the 13 Gyr old population in the halo. The final result is a birth population of 170 million binary systems within 2 kpc.
@@
For binaries where both the primary and secondary have reached the end of the post-AGB phase, we assume those are two white dwarfs that will not produce a Gaia-detectable binary as both components are too faint. For binaries where only the primary has reached the post-AGB phase, we assume these are white dwarf + star binaries. Following the prescription of El-Badry et al. (2024); Yamaguchi et al. (2024), we assume that 10% of these white dwarf + star binaries with initial separations between $2-6$ AU form wide post-common envelope binaries with separations that are half their initial separation. The white dwarfs are assumed to be dark and their masses are determined from their zero-age main sequence mass using the initial-final mass relation of Kalirai et al. (2008). The remaining white dwarf + star binaries are assumed to either have orbital periods too short ($\lesssim 10$ days) or too long ($\gtrsim 10^{4}$ days) that they could not be detected astrometrically by Gaia in DR3. We also remove binaries where either star is currently or has previously overflowed its Roche lobe, under the assumption that the periods will be too short to detect astrometrically in Gaia DR3 due to orbital shrinking.
@@
where $l$ is the luminosity ratio $l=L_{2}/L_{1}$ of the binary system and $q_{l}$ is the mass ratio $q_{l}=M_{2}/M_{1}$ (van de Kamp, 1975; Belokurov et al., 2020). By definition here, the primary (component 1) is brighter, but not necessarily more massive. The mass ratio $q$ is defined to be $0\leq q\leq 1$. Thus, $q_{l}=q$ if component 1 is both brighter and more massive, otherwise $q_{l}=1/q$. Note that the photocenter orbit semimajor axis is maximized for dark companions (i.e., $l=0$) and large mass ratios (i.e., $q_{l}\gg 1$), and is zero for equal-mass companions of the same luminosity ($l=1$, $q_{l}=1$). Finally, we note that all these luminosity-dependent quantities are specifically in the Gaia G-band.

### v2

@@ Abstract
In June 2022, the Gaia mission released a catalog of astrometric orbital solutions for 168,065 binary systems, by far the largest such catalog to date. The catalog’s selection function is difficult to characterize because of choices made in its construction. Understanding the catalog’s selection function is required to model and interpret its contents. We use a combination of analytic and empirical prescriptions to construct a function that computes the probability that a binary with a given set of properties would have been published in the Gaia Data Release 3 astrometric orbit catalog. This is a complementary approach to the more accurate but significantly more computationally expensive approach of El-Badry et al. (2024). We also construct a binary population synthesis model to validate our characterization of the selection function, finding good agreement with the actual Gaia NSS catalog, with the exception of the orbital eccentricity distribution. The NSS catalog suggests high-eccentricity orbits are relatively uncommon at intermediate periods $100\lesssim P_{orb}\lesssim 1000$ days. As an example application of the selection function, we estimate the Gaia DR3 detection probabilities of the star + BH binaries Gaia BH1 and BH2, and find them to be 0.38 and 0.27, respectively. Compared to the values obtained by detailed modeling in El-Badry et al. (2024), the probabilities are identical for BH1, and within a factor of 2 for BH2. We also estimate the population of Sun-like star + BH binaries in the Galaxy to be $\sim 3000$ for $400100<P_{orb}<$ 400 day, $<800$ for $1000400<P_{orb}<$ 1000 day, and $<12,000$ for $15001000<P_{orb}<$ 1500 day.
facilities: Gaiasoftware: isochrones (Morton, 2015), galaxia (Sharma et al., 2011), astropy (Astropy Collaboration et al., 2013, 2018, 2022), Matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011), SciPy (Virtanen et al., 2019), healpy (Zonca et al., 2019), mwdust (Bovy et al., 2019), COSMIC (Breivik et al., 2020)
@@ 1 Introduction
Gaia is an astrometric space mission surveying over 2 billion stars in the Milky Way (Gaia Collaboration et al., 2016a). Its main mission operated from July 2014 through December 2019, and its extended mission operated through January 2025. The data products and catalogs generated from the photometry, astrometry, and spectroscopy from Gaia are shared publicly in data releases that occur every several years.
@@
In this paper, we perform an empirical forward model of the Gaia NSS astrometric orbits catalog. We focus on the solutions that have an astrometric orbit (as compared to those that only have a spectroscopic orbit) as our primary goal is to understand the selection function of stellar-mass black hole + star binaries, which have thus far only been found in the astrometric orbits catalog. However, we emphasize that the selection function is applicable to any type of binary system, e.g., star + star, or star + planet. The selection function is defined in terms of observables (e.g., the size of the photocenter’s orbit; Equation 1). Although different companion luminosities and masses will result in different observables and thus different detection probabilities, Gaia is agnostic to whether the secondary component of an unresolved binary is a star, planet, or black hole. The purpose of the selection function is to quantify how different those detection probabilities are.
Our forward modeling approach combines analytic models and results from the individual-epoch forward modeling of El-Badry et al. (2024) to produce fast, empirical functions to calculate the probability of a source being included in the NSS astrometric orbits catalog.
The rest of this paper is organized as follows. In §2 we review the processing pipeline of the Gaia NSS astrometric orbit catalog and identify the most important steps to model. In §3 we describe the binary population synthesis model that we compare to the NSS catalog. In §4 we forward model the NSS astrometric processing cascade, and the results are presented in §5. Applications, strengths and shortcomings, and possible extensions of the forward model are discussed in §6. We provide a final summary and conclusions in §7.
@@ 2 Gaia astrometric non-single star processing and catalog
The astrometric NSS processing and resultant catalog were designed with the intention of prioritizing accurate binary parameters over completeness. This resulted in a series of many filtering steps to remove potentially spurious solutions, at the expense of many accurate solutions. Our goal is to understand the selection effects of this filtering process, as well as determine which filters are most important to model. The astrometric NSS processing described in Halbwachs et al. (2023) is summarized below, and the portions most relevant to the forward modeling are highlighted in Figure 1; see also Figure 1 of Halbwachs et al. (2023) for a summary.
Figure 1: Summary of the NSS astrometric orbit cascade, illustrating the criteria that must be satisfied for a source to be included in the NSS astrometric orbit catalog.
@@ 2.1 Astrometric orbit solutions
The Gaia DR3 NSS catalog is divided into four tables depending on the type of NSS model used; each table is further subdivided by the particular sub-model used (see §2.1. “Table contents” of Gaia Collaboration et al. (2023b)). This paper is only concerned with modeling the subset of solutions that have astrometric orbits, i.e. Orbital and AstroSpectroSB1 solutions from the nss_two_body_orbit table. We collectively refer to these as the NSS astrometric orbit solutions. Unless otherwise stated, we will also just use the shorthand “NSS solutions” or “NSS catalog” to describe this dataset, since we are only concerned with astrometric models in this paper.
@@ 2.2 Selection of stars to be processed
In the astrometric NSS processing, an initial filtering was performed on the entire Gaia DR3 gaia_source  catalog to define the input catalog, a subset of sources that were good NSS candidates. Starting from $1.5\times 10^{9}$ sources with single-star astrometric solutions, several filtering steps were implemented to produce the input catalog of $4.1\times 10^{6}$ stars. Only stars in this input catalog were eligible to be fit with an astrometric NSS model. As described in Halbwachs et al. (2023) §2.1, the filters are on the source’s apparent magnitude, goodness of fit of a single star solution, number of visibility periods (i.e., independent observations), [^1] Using the number of visibility periods provides a more reliable estimate for position and proper motion uncertainties than the total number of good along-scan observations (Lindegren et al., 2018). image parameter determination metrics, and color excess. The filters on the image parameter determination metrics and color excess are designed to remove marginally resolved wide binaries. Gaia DR3 has an effective angular resolution of around 0.7 arcsec [^2] Although Gaia DR3 has an angular resolution of 0.18 arcsec (Lindegren et al., 2021), it is limited spatially by crowding, making its effective angular resolution 0.7 arcsec.; at smaller separations the completeness of close source pairs drops (Gaia Collaboration et al., 2021a).
In our forward model, we will only model the filter on the goodness of fit of a single star solution ($\texttt{ruwe}>1.4$, §4.2). The filters on apparent magnitude and visibility periods are accounted for in the ruwe  filter. As ruwe  is a measurement of the significance of deviations from a single-star model, sources that are faint or have too few visibility periods will have astrometric uncertainties too large to have a significant deviation. In addition, our forward model does not explicitly model the image parameter determination metrics or color excess, because the wide binaries removed by these sources will be removed by the $\texttt{ruwe}>1.4$ filter. Using a binary population synthesis model (§3), we find this angular resolution corresponds to binaries within 2 kpc having orbital periods several centuries to millennia. These orbital periods are much too long to receive an astrometric orbital solution; they have $\texttt{ruwe}>1.4$ because their marginally resolved nature leads to spurious epoch astrometry. See also Tokovinin (2023) and Appendix B of El-Badry et al. (2024).
2.3 Model cascade and main processing
Subsequently in the astrometric NSS processing pipeline, the $4.1\times 10^{6}$ stars from the input catalog were fit with a variety of astrometric NSS models. The models were tried in the following order [^3] We do not model the “variability-induced movers” branch in the processing cascade; this describes sources with astrometric motion in unresolved binaries where one source is photometrically variable. We do not consider variable sources in our catalog so this is not relevant.:
@@ 2.3 Model cascade and main processing
  • orbital (Orbital solution, nss_two_body_orbit table; 12 parameters44410 parameters were used for a circular or psuedo-circular orbit (see §5.2 “Pseudo-circular and circular orbital solutions” of Halbwachs et al. (2023)). 15 parameters were used if the orbital solution was combined with a spectroscopic solution (AstroSpectroSB1 solution). The astrometry-only and spectroscopic-only inferred parameters were only combined if they matched (see §5.3. “Selection of the orbital solutions” of Halbwachs et al. (2023)).).
In the NSS astrometric processing pipeline, there are two steps: main processing and post-processing. If a source passes a given model’s main processing criteria, it moves into the post-processing branch of that model; if a source does not pass a given model’s main processing criteria, the next model in the cascade is tried. If a source passes both a given model’s main processing and post-processing criteria, it is accepted as a solution and is included in that model’s catalog. If a source does not pass the post-processing criteria, it is removed and given a single star solution.
For each model, the criteria for main processing acceptance were based on the the significance of the model and the significance of the parallax measurement. The specific criteria are summarized in Table 1 of Halbwachs et al. (2023). If an acceptable solution was found, subsequent models were not tried. Note that the acceleration models precede the orbital model. Thus, if an acceptable acceleration model was found, an orbital model was not tried, even if an orbital solution would have led to a better fit.
In our forward model, we model the variable acceleration (referred to as 9-parameter acceleration), constant acceleration (referred to as the 7-parameter acceleration), and the orbital components of the astrometric cascade. For each of these three models, we model the main processing filters on the significance of the model and the parallax measurement.
Throughout this paper we discuss acceptance of acceleration and orbital models. When we say an acceleration model is accepted or rejected, we are specifically referring to the main processing step unless otherwise stated. This is because only the main processing branch of the acceleration models is relevant to the orbital solutions. If an acceleration model is rejected from the main processing, it is a candidate for an orbital solution. If an acceleration model is accepted in the main processing, it will not be a candidate for an orbital solution regardless of whether it is accepted or not in the subsequent acceleration post-processing. In contrast, when we say an orbital model is accepted or rejected, we are specifically referring to acceptance in both the main and post-processing steps. This is because we are interested in the final astrometric orbits catalog, which requires passing the post-processing criteria.
@@ 2.4 Orbital solution model acceptance
For sources in the NSS astrometric processing pipeline where an orbital model fit was attempted in the main processing, there are three additional conditions determining whether the solution was accepted in the post-processing and included in the final orbital solution catalog: a criterion on the goodness-of-fit $(F_{2})$, a second more stringent filter on the significance of the model, and a filter on the error of the eccentricity measurement $\sigma_{e}$ (Halbwachs et al., 2023).
In our forward model, we do not model the filter on $\sigma_{e}$, as it removes no more than a few percent of solutions and does not bias the inferred distributions of orbital parameters. Detailed justification is provided in Appendix A. Visual inspection of Figure 3 of Halbwachs et al. (2023) also suggests the filter on $\sigma_{e}$ has minimal impact on the final constructed catalog.
In our forward model, we also do not model the filtering on $F_{2}$ because we assume that good orbital solutions will have good fits. $F_{2}$ is expected to follow a normal distribution with mean 0 and standard deviation 1; even if the uncertainties are poorly characterized, $F_{2}<25$ is still a 25-$\sigma$ threshold. It may also allow false positives to pass through, but is unlikely to remove real solutions that have passed the other filtering steps (with the exception of binaries in hierarchical triple systems). We provide further justification of this choice in Appendix B.
@@ 2.5 Final NSS astrometric orbit catalog
Figure 2: 1-D and 2-D PDFs in the period $P_{orb}$, parallax $\varpi$, eccentricity $e$, photocenter orbit semimajor axis $a_{0}$, ruwe, apparent magnitude $G$, and parallax over parallax error $\varpi/\sigma_{\varpi}$ of the actual NSS astrometric orbits catalog (black) as compared to forward model (red, described in §35). In general, the forward modeling does a good to reasonable job in matching the actual NSS distributions including the correlations, with the exception of the eccentricity distribution.
The black outline in Figure 2 shows the distribution of orbital period, parallax, eccentricity, photocenter orbit semimajor axis, renormalized unit weight error, apparent magnitude, and parallax over parallax error for the final set of 168,065 Gaia DR3 NSS astrometric orbit solutions. Some of the correlations are clearly due to the filtering procedure (e.g. orbital period vs. parallax over parallax error).
@@ 3 Binary population synthesis
We emphasize that the population synthesis is used primarily to validate the output of the selection function, and that it is not used in determining the selection function itself. Once the selection function has been determined, different population synthesis models can be explored to determine whether they produce different predicted numbers of detectable black holes (e.g., Nagarajan et al., 2025), although such a study is out of the scope of the work presented here.
@@ 3.2 Stellar population model (binary properties and evolution, IMF, synthetic photometry)
To forward model the NSS catalog, we cannot directly use the stellar population generated by galaxia since it consists of only single stars. Hence, we combine galaxia positions, metallicities, and ages with a binary star population based on Moe & Di Stefano (2017), which is a joint probability distribution in primary mass $M_{1}$, mass ratio $q=M_{2}/M_{1}$, orbital period, and eccentricity. Moe & Di Stefano (2017) infer the distribution over $0.8<M_{1}<40M_{\odot}$, $10.1<q<$ 1, $80.15<\log P_{orb}/\textrm{day}<$ 8, and $10<e<$ 1. For reference, about 60% of Sun-like stars are single, while this decreases to nearly 0% for OB-type stars.
@@
We assume that the mass distribution of primaries and singles are drawn from a Kroupa initial mass function (Kroupa, 2001). This results in a multiplicity frequency [^6] Following the definition of Reipurth & Zinnecker (1993), the number of binary systems divided by the number of binary systems and single stars. of 0.21 and a companion probability [^7] Again following Reipurth & Zinnecker (1993), the number of stars in binary systems divided by the total number of stars. of 0.35. This multiplicity frequency might seem significantly lower than binary frequency of $30-40\%$ typically quoted for solar-type stars (e.g., Raghavan et al., 2010; Moe & Di Stefano, 2017). This is because the binary statistics in Moe & Di Stefano (2017) only extend down to $0.8M_{\odot}$; to extrapolate the multiplicity statistics down to $0.08M_{\odot}$, COSMIC assumes the binary fraction decreases linearly from $\approx 40$% at $0.8M_{\odot}$ down to 0 at $0.08M_{\odot}$. This choice of implementation is somewhat arbitrary (given that M-dwarfs appear to have a binary frequency $\sim 20\%$; e.g., Laithwaite & Warren 2020), but ultimately most of these sources are too faint to be included in the astrometric orbits catalog, so the model is not sensitive to the particular binary fraction assumed. Based on an estimate using the extinction-corrected absolute magnitudes of the NSS sources, we find that main sequence M-dwarf binaries comprise $\lesssim 5\%$ of the NSS catalog. Similarly, we do not simulate substellar companions, as they are a negligible part of the NSS catalog; Gaia Collaboration et al. (2023b) find that only 1% of NSS companions are brown dwarfs or exoplanets.
Assuming that the majority of binary systems are unresolved [^8] Using the criterion for resolvability given by Equation (B1) in El-Badry et al. (2024) and assuming the binary separation is equal to the semimajor axis, we find that 80% of binary star systems will be unresolved., one galaxia star corresponds to one COSMIC stellar system. This corresponds to 150 million binary systems within 2 kpc at the present day. In order to properly generate the birth population, the stellar progenitors of the compact objects must be accounted for [^9] Note that compact objects are not included in galaxia.. We use the MIST stellar evolution models (Dotter, 2016; Choi et al., 2016) to evolve a population of stars to the same age distribution of the thin disk, thick disk, stellar halo, and bulge. We then determine how many of the primary stars have or will imminently become compact objects (defined as carbon burning for massive stars, and the end of the post-AGB phase for low/intermediate mass stars), and apply a correction factor to the number of stars that need to be generated. This correction is 7.5% for the 5–7 Gyr old population in the thin disk, 11.5% for the 11 Gyr old population in the thick disk, and 12.5% for the 13 Gyr old population in the halo. The final result is a birth population of 170 million binary systems within 2 kpc.
@@
For binaries where both the primary and secondary have reached the end of the post-AGB phase, we assume those are two white dwarfs that will not produce a Gaia-detectable binary as both components are too faint. For binaries where only the primary has reached the post-AGB phase, we assume these are white dwarf + star binaries. Following the prescription of El-Badry et al. (2024); Yamaguchi et al. (2024), we assume that 10% of these white dwarf + star binaries with initial separations between $2-6$ AU form wide post-common envelope binaries with separations that are half their initial separation. The white dwarfs are assumed to be dark and their masses are determined from their zero-age main sequence mass using the initial-final mass relation of Kalirai et al. (2008). The remaining white dwarf + star binaries are assumed to either have orbital periods too short ($\lesssim 10$ days) or too long ($\gtrsim 10^{4}$ days) to be detected astrometrically by Gaia in DR3. We also remove binaries where either star is currently or has previously overflowed its Roche lobe, under the assumption that the periods will be too short to detect astrometrically in Gaia DR3 due to orbital shrinking.
@@
where $l$ is the luminosity ratio $l=L_{2}/L_{1}$ of the binary system and $q_{l}$ is the mass ratio $q_{l}=M_{2}/M_{1}$ (van de Kamp, 1975; Belokurov et al., 2020). By definition here, the primary (component 1) is brighter, but not necessarily more massive. The mass ratio $q$ is defined to be $0\leq q\leq 1$. Thus, $q_{l}=q$ if component 1 is both brighter and more massive, otherwise $q_{l}=1/q$. The photocenter orbit semimajor axis is maximized for dark companions (i.e., $l=0$) and large mass ratios (i.e., $q_{l}\gg 1$), and is zero for equal-mass companions of the same luminosity ($l=1$, $q_{l}=1$). We note that all these luminosity-dependent quantities are specifically in the Gaia G-band.
3.3 Gaia-specific observational quantities
Finally, the population synthesis model also requires information about Gaia’s observational properties: the number of observations and the measurement error. To obtain an estimate for the number of visibility periods $N_{vis}$ in the population synthesis, we use the Gaia Observation Forecast Tool (GOST) [^10] https://gaia.esac.esa.int/gost/, which predicts when any given part of the sky crosses Gaia’s focal plane. To calculate $N_{vis}$, we evaluate the number of predicted scans at 49152 evenly spaced points on the sky, discard a random 10% of the predictions to mimic losses due to dead time [^11] The GOST page suggests 20% would be more appropriate, but we find 10% gives good agreement with the Gaia DR3 catalog., then calculate how many of the remaining predicted observations are separated by at least 4 days.
We adopt the along-scan per-CCD centroiding error $\sigma_{AL}(G)$ from Figure 3 of Holl et al. (2023). $\sigma_{AL}(G)$ is a function of the source magnitude $G$. As described in Lindegren (2018), for $12\lesssim G\lesssim 17$ the formal centroiding precision is limited by photon noise (leading to Poisson errors), for $G\gtrsim 17$ background noise dominates and the error rises more quickly than Poisson, and for $G\lesssim 12$ the CCD gates limit the number of photons and the centroiding precision is independent of magnitude. The $\sigma_{AL}(G)$ we adopt from Holl et al. (2023) is empirically determined from Gaia DR3 data, and is significantly larger ($>0.3$ dex) than the formal error for bright ($G\gtrsim 12)$ sources.